Hi I need help in a certain problem on my algebra homework. "Four times the larger of 2 numbers is equal to seven times the smaller. The sum of the numbers is 22. Find the numbers. Thank you Mr./Ms. Tutor :)
Answer by josgarithmetic(39617)

EDIT: Parts of this answer seemed to be missing.
You describe this: numbers a,b.
a 4b=7a
and
a+b=22.

Just work with the equations and you'll find the numbers.
b=22-a
4(b)=7a
4(22-a)=7a
88-4a=7a
88=11a
8=a

From that, find b.
b=22-a=22-8=14